The input voltage to an audio amplifier is 25mV. An oscilloscope connected across a 1MΩ resistor in series with the input to the amplifier indicates 1V peak-to-peak. If the voltage across a 10KΩ load is 3.5V. Determine :
a) the current, voltage and power gains
b) the amplifier's input resistance
1
Expert's answer
2020-10-16T10:59:55-0400

Explanations & Calculations


  • Vrms value of that applied across the "\small 1M\Omega" resistor = "\large \frac{V_{p-p}}{2\sqrt2} \small = 0.3536V"
  • Therefore, the input current = "\large \frac{0.3536V}{1M\Omega} \small= 0.3536\mu A"
  • Current drawn by/ supplied to the load = "\large \frac{3.5V}{10k\Omega} \small = 0.35mA"


  • Therefore, the current gain = "\large\frac{I_{out}}{I_{in}} \small=\large\frac{0.35mA}{0.3536\mu A} \small = \large \frac{350\mu A}{0.3536\mu A} \small = \small\bold{989.82}"
  • Voltage gain = "\frac{V_{out}}{V_{in}}= \frac{3.5V}{25mV} = \frac{3500mV}{25mV} = \small \bold{140}"


  • Power gain = Voltage gain "\small \times" current gain = "\small 140 \times 989.82 = \bold{1.39\times 10^5}"


  • Input resistance = "\large \frac{\text{input voltage}}{\text{input current}} \small = \large \frac{25mV}{0.3536\mu A} \small = \bold{70.701k\Omega}"